OpenStudy (jthec):

Glucose is pumped into some animal cells by a symporter powered by the simultaneous entry of Na+. The entry of Na+ provides a free-energy input of 10.8kJ/mol under typical cellular conditions (extracellular [Na+] =143mM, internal [Na+] =14mM, and membrane potential = 50mV). How large a concentration of glucose can be generated by this free-energy input ( what is the ratio of glucose in / ratio of glucose out)

OpenStudy (agreene):

I feel like maybe they're giving you a bunch of useless information (word questions generally do). I think: extracellular [Na+] = 143mM and cellular [Na+] = 14mM thus ratio of cellular : extracellular would be \[\frac{14}{143} \]

OpenStudy (agreene):

However, if you have some \(\Delta G\) function for this particular example; you might need it.

OpenStudy (agreene):

I just realised they gave you a mV ... so they're expecting you to use one of those thermodynamic potential equations that I dont remember

OpenStudy (jthec):

I used the equation Delta G = RTln(C2/C1) +ZF Delta psi

OpenStudy (jthec):

Then plugged the DG back into the equation Delta G = RTln(C2/C1) to solve for C2/C1

OpenStudy (jthec):

I know the answer is 66 (concentration inside to outside) and I got the answer doing it that way, but did I do it wrong?